OpenStudy (anonymous):

i really need help. watz the answerd to this formula. D=1/5fk for k

OpenStudy (anonymous):

your question is not clear oswaldo

OpenStudy (stormfire1):

use parenthesis so it's more clear

OpenStudy (anonymous):

well it say find the varible. for this formula D=1/5fk for k

OpenStudy (stormfire1):

The problem is that it's unclear whether it's D = (1/5f)k D = (1/5fk) D = (1/5)fk And it matters which way it is

OpenStudy (anonymous):

in the problem is say k=

OpenStudy (stormfire1):

From above, you could get 3 different answers. We need to know whether f and/or k are supposed to be part of the denominator of the fraction.

jhonyy9 (jhonyy9):

D=1/5fk 5D=fk 5D/f =k

jhonyy9 (jhonyy9):

this is ok ?

OpenStudy (anonymous):

solve the formula for the specified variable D=1/5fk for k k=

OpenStudy (anonymous):

thats how the problem looks

jhonyy9 (jhonyy9):

so ok but fk is denominator or numerator ?

OpenStudy (stormfire1):

The teacher should be shot if something as simple as this isn't 100% clear. On the paper, does it look like:\[D=\frac{1}{5}fk\]If so, jhonyy9's answer is correct.

jhonyy9 (jhonyy9):

thank you stormfire1

OpenStudy (anonymous):

yes it looks like that. nd thank u for helping mi

OpenStudy (stormfire1):

np

jhonyy9 (jhonyy9):

my pleasure

jhonyy9 (jhonyy9):

bye